5-2(y-5)=27 how do you solve and check that and could you show all steps and give a free ex. please and thank you
Answer by ankor@dixie-net.com(22740)

: 5-2(y-5)=27 how do you solve and check that and could you show all steps
:
: 5 - 2(y-5) = 27
:
First multiply what is inside the brackets, remember if there is a negative on the outside of the brackets, you change the signs (+/-) inside the brackets:
:
5 - 2y + 10 = 27
:
-2y + 5 + 10 = 27
-2y + 15 = 27
:
Subtract 15 from both sides:
-2y + 15 - 15 = 27 - 15
:
-2y = 12
:
Divide both sides by -2, this makes it +1y or just y,
:
y = 12/-2
:
y = -6
:
:
Prove that y = -6 is correct, substitute -6 for y in the original equation:
: 5 - 2(y-5) = 27
:
5 - 2[-6 - 5] = 27
:
5 - 2[-11] = 27; a minus times a minus is a plus
:
5 + 22 = 27; equality proves our solution
:
How about this? Did it help you on your way to understanding this?